OpenStudy (anonymous):

What is the approximate distance from (-4, 3) to (1, 0)?

OpenStudy (saifoo.khan):

5.83

OpenStudy (anonymous):

\[\sqrt{(0-3)^2+(1-(-4)^2}=\sqrt{9+25}=\sqrt{37}\]

OpenStudy (anonymous):

sorry \[\sqrt{34}=5.830\]not square 37